OpenStudy (anonymous):

FAN and MEDAL. Choose the point-slope form of the equation below that represents the line that passes through the points (-3, 2) and (2, 1). y + 3 = -5(x - 2) y - 2 = -5(x + 3) y + 3 = -one fifth(x - 2) y - 2 = -one fifth(x + 3)

OpenStudy (anonymous):

okay.... so do you know what point-slope form is?

OpenStudy (studygurl14):

find slope, then plug known values into point-slope form

OpenStudy (anonymous):

point -slope form is y1 - y2 = m(x1 - x2)

OpenStudy (anonymous):

your slope is 1/5

OpenStudy (anonymous):

so, this narrows down your answer to either c or d

OpenStudy (anonymous):

-1/5

OpenStudy (anonymous):

oh, yeah... -1/5 sorry :)

OpenStudy (anonymous):

So you use one of the points and plug it into the equation to get your answer. In this problem, they used (-3, 2)... SO, your answer is ???

OpenStudy (anonymous):

no clue thats why i posted it on here

OpenStudy (anonymous):

okay, so you know that the slope is -1/5, right? so, they used the point (-3, 2) in this problem and plugged it into the equation, and the coordinates are x, y, right?

OpenStudy (anonymous):

so, you would use y1 - y2 = m(x1 - x2) plug the numbers in and you get: y + 3 = -1/5(x - 2)

OpenStudy (anonymous):

that is which answer? C or D?

OpenStudy (anonymous):

it is c. hope I could help. :) good job.... normally most people don't cooperate. they just sit there and wait for me to answer the question. :)

OpenStudy (anonymous):

tbh thats really all i was doing , thank you for being so smart i wouldnt have found the answer by myself